# Surf-NeRF: Surface Regularised Neural Radiance Fields

### v1

@@ Abstract
Neural Radiance Fields (NeRFs) provide a high fidelity, continuous scene representation that can realistically represent complex behaviour of light. Despite recent works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited. We demonstrate how curriculum learning of a surface light field model helps a NeRF converge towards a more geometrically accurate scene representation. We introduce four additional regularisation terms to impose geometric smoothness, consistency of normals and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models. Our approach yields improvements of 14.4% to normals on positionally encoded NeRFs and 9.2% on grid-based models compared to current reflection-based NeRF variants. This includes a separated view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene. We demonstrate compatibility of our method with existing NeRF variants, as a key step in enabling radiance-based representations for geometry critical applications. Project page: https://roboticimaging.org/Projects/SurfNeRF
@@ 1 Introduction
Figure 1: Surf-NeRF uses the properties of surface light fields to regularise a NeRF by querying the local region of samples. These samples are used to improve consistency and smoothness in these regions accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([36], left) whilst maintaining visual fidelity. By changing how frequently regularisation occurs during training, we can control the strength of the effect.
Whilst re-parameterisation of the radiance in the scene as in Ref-NeRF [36] has shown improvements over previous state-of-the-art variants, there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density. It is still common for the NeRF to place regions of density behind surfaces or in front of the camera as detached geometry (floaters) to explain complex phenomena. This leads to a non-realistic geometric scene representation, as in Figure 1.
Applications like robotics and 3D modelling require both accurate geometry and realistic appearance. The poor physical geometry currently poses a large hurdle to the widespread adoption of NeRFs as a scene representation for geometry critical tasks. Robotic manipulation and autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows. Similarly, traditional structure-from-motion pipelines have great difficulty in reconstructing visually complex objects including reflective and transparent surfaces.
We address geometric inaccuracy arising from view-dependent phenomena with the insight that there are multiple formulations of the plenoptic function which can produce viable scene representations. A surface light field [41] describes the plenoptic function as light originating from a geometrically smooth surface. Our insight lies in that we can push the NeRF towards a more geometrically accurate representation in the same rendering framework using additional geometric and appearance based regularisation via curriculum learning. We use a first surface assumption to describe the location of geometry in the scene. Using a second sampling of the NeRF at these points, we regularise density to produce smoothly-varying normals and thin, continuous sheets of density which more realistically represent scene geometry. By enforcing view-dependent properties of a reflection light model at these points, we enable a separation of the Lambertian appearance of a scene from the view-dependent component, reducing shape-radiance ambiguity and encouraging more correct structure of the surface.
In this work, we make the following contributions:
@@
    We leverage curriculum learning of a NeRF towards a more accurate geometric scene representation which maintains visual fidelity whilst refining the density representation of the scene.
Whilst we benchmark our approach on state-of-the-art physics based NeRF variants, our methodology may also be applied to other NeRF frameworks.
@@ 2 Related Work
Neural Scene Representations: Neural field approaches [35, 14] to scene representation produce continuous and often high visual fidelity depictions which are able to be queried anywhere within the training set, balancing visual and geometric fidelity. In this work, we seek to improve the geometry of a NeRF whilst maintaining visual fidelity.
@@
NeRFs [27] leverage the efficiency of ray-based construction similar to light fields with a volumetric scene representation of points which emit light. Subsequent works have improved the fidelity of representation [3, 4, 6], recovery of a static scene [32, 34, 26] and its performance around view-dependent phenomena such as reflections [36, 22]. However, relying on a purely volumetric rendering approach still allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [31, 23, 29] or where light does not follow the physical model of a straight ray through the scene [8]. Additional regularisation terms have been shown to dramatically improve the quality of rendering in these cases [29, 23]. Improving the accuracy of normals in the scene also significantly helps learning around complex appearance [25, 36]. More recent works [37] have shown that ray-tracing multiple reflection rays can improve the representation of reflections, albeit having extremely high computation cost. In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries. This approach explains complex visual phenomena entirely through view-dependent appearance, whilst maintaining the formulation of a single-pass volumetric rendering approach and its quality. We achieve this using a surface light field model and local characteristics of geometry and appearance of the scene.
@@
Surface Representations: Signed distance fields (SDFs) have garnered significant attention [44, 45, 16, 18, 2, 30] as they can represent smooth and continuous geometries in space and may be generated from multi-view constraints alone. Applying view-dependent colour channels to the SDF [38, 44, 18, 39, 40] in a similar fashion to NeRFs has demonstrated improved accuracy and continuity of representation. For diffuse objects, this representation provides a smooth and high quality reconstruction, however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented. Some works have included a reflection parameterisation [21], however struggle to reconstruct 3D scenes with fine detail, thin structures and sharp changes compared to volumetric scenes. Appearance and geometry have an intrinsic relationship within neural representations, with degradation in geometric accuracy often resulting in altered geometry to meet appearance [18]. Our work maintains a volumetric scene representation allowing for complex geometry, but with separated view-dependent and -independent appearance. Given NeRFs are a high fidelity visual representation, we are concerned only with introducing a surface-like structure to the density field through volumetric rendering. This improves the geometry and consistency of novel view rendering by considering the NeRF as learning density constrained to a surface with smooth view-dependent terms.
@@
Appearance Vs. Geometry: Inverse rendering seeks to produce a definite [20] separation of the scenes appearance, geometry and environment, which are necessary to create realistic renderings of scene’s under new conditions. Performing this without prior knowledge of the scene proves to be an immensely difficult task [46, 11, 15, 43], however, utilising physically-based rendering fused with learnt appearances [10, 12] provides a sufficiently constrained framework to acquire the components of the scene. Radiance field approaches seeking to learn an accurate scene representation under few view scenarios have used depth consistency [33], additional depth supervision [49] and regularisation using priors [29]. Incorporating an understanding of the physical interactions of light within a scene and its effect on appearance [47] under a solid surface assumption enables accurate geometry to be learnt alongside appearance. In the presence of specularities and other visual phenomena it is difficult to disentangle where appearance has been baked into geometry [42]. Without the need for recovering a bidirectional reflectance distribution function (BRDF), light field approaches [19] provide a more generalised framework. This enables a clear separation between appearance and geometry, by reducing the need to acquire environmental view-dependent effects. Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric rendering framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
@@ 3 Method
Figure 2: An overview of the Surf-NeRF methodology. We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene. Sampling at this point in multiple directions, and at points nearby using points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region. We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by sampling points through multiple directions. Using neighbouring points we also regularise geometric smoothness on density $\tau$ and consistency of normals $\hat{\mathbf{n}}$ leading to improved, continuous geometry.
@@ 3.1 Preliminaries
Few real surfaces have an entirely view-dependent appearance. Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance, $S$, and specular (or more generally a view-dependent) component, $R$ [41, 24], $L(\mathbf{u},\boldsymbol{\omega})=S(\mathbf{u})+R(\mathbf{u},\boldsymbol{\omega})$. These intrinsic decompositions are defined for a point on the surface, $\mathbf{u}$, and viewing direction, $\boldsymbol{\omega}$. The diffuse reflectance varies only with position over the surface, whilst the view-dependent component captures elements like reflection and refraction at the surface [20].
A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a continuous geometry in the scene. This accordingly models phenomena such as volumetric scattering, reflection or transmission as a function of surface position rather than in a volumetric (NeRF) or physics-based (rendering engine) manner. In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance looking along a ray is maintained [41]. Our approach is motivated by this decoupling to encourage Lambertian radiance to exist on smooth sheets of density, or surfaces, with a view-dependent colour.
@@ 3.2 Model
Our proposed method builds on the reflection parameterisation in Ref-NeRF [36], which splits the scene into a diffuse and specular appearance term similar to a surface light field. This parameterisation provides enhanced results across Lambertian and specular scenes over original NeRF variants [27, 3] by learning a spatially varying diffuse colour $\mathbf{c}_{d}$, specular tint $\mathbf{s}$, rendering normal $\hat{\mathbf{n}}^{\prime}$ and a view-dependent specular colour $\mathbf{c}_{s}$ for each point in the scene. The final colour of a ray in this parameterisation is given as $\mathbf{c}=\mathbf{c}_{d}+\mathbf{s}\mathbf{c}_{s}$, a linear combination of the diffuse and specular terms. Ref-NeRF struggles to entirely separate Lambertian and specular appearance, utilising density instead to explain complex phenomena such as non-planar reflection, anisotropic reflection and interreflection, leading to the results seen in Figure 1. We apply our regularisation losses denoted by $\mathcal{L}$ to ZipNeRF [6], leveraging its state of the art performance with grid-based encoding [28], using the Ref-NeRF parameterisation. We make no modifications to the model itself beyond this, but include a second, data-driven sampling to impose physically-inspired regularisation. We maintain two proposal networks and one NeRF network with 64 and 32 samples per ray. Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage, as we show in Section 4.
@@ 3.3 Sampling Radiance at a Surface
We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray insects a surface in the scene, as shown in Figure 3. We formulate local regularisation terms based on surface light field properties and the current scene geometry and appearance at this location.
@@
We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by dense points closest to the camera. Our candidate surface is the first point along a ray with weight $w(\mathbf{x}_{\star})$ greater than the median weight of the ray $\mathrm{med}({\mathbf{w}}(\mathbf{x}))$, preventing regularisation from occurring behind the true location of the surface. This minimises sampling around points which are occluded and therefore not well positioned with multi-view constraints. Choosing the median ensures that this selection is more robust to skewed weight distributions, particularly early during training. This is shown in Figure 2 left.
Using the point ($\mathbf{x}_{\star}$), origin ($\mathbf{o}$), direction ($\mathbf{d}$), surface normal ($\mathbf{\hat{n}}_{\star}$), and covariance ($\mathbf{\Sigma}_{\star}$) of this sample, we generate two new batches; a spatial batch to regularise density and a directional batch to regularise appearance.
@@
To produce virtual rays which can be cast during training, we use our sphere samples to define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta. The origins and covariances for these rays,
@@
We use these new virtual rays to produce two batches, querying the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the distribution of view-dependent colours $\mathbf{c}_{s}$ through a range of viewing angles. We refer to these as the spatial and directional batches. Below we outline the locations $\mathbf{x}$, and directions $\mathbf{d}$ through which we sample the NeRF in these batches. We also visually depict our sampling schemes in Figure 3.
@@
$\quantity{\mathbf{x}_{\phi},\,\mathbf{d}_{\phi}}=\quantity{\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}},$ (2)
@@
where $\sigma_{\star,r}$ is the radial variance of the sample as per mip-NeRF [3]. This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail compared to those further away. Importantly, our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training. We characterise this behaviour in the supplement. As we do not care about the colour of these samples off the surface, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.

### v2

@@ Abstract
Neural Radiance Fields (NeRFs) provide a high fidelity, continuous scene representation that can realistically represent complex behaviour of light. Despite works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited. We demonstrate how both curriculum learning of a surface light field model and using a lattice-based hash encoding helps a NeRF converge towards a more geometrically accurate scene representation. We introduce four regularisation terms to impose geometric smoothness, consistency of normals, and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models. Our approach yields 28% more accurate normals than traditional grid-based NeRF variants with reflection parameterisation. Our approach more accurately separates view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene. We demonstrate compatibility of our method with existing NeRF variants, as a key step in enabling radiance-based representations for geometry critical applications.
@@ 1 Introduction
Figure 1: Surf-NeRF uses a permutohedral lattice encoding and the properties of surface light fields to regularise local regions of a NeRF. This improves consistency and smoothness in these regions by accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([43], left). Changing the frequency of regularisation balances between geometric fidelity and appearance separation.
Re-parameterisation of the radiance in the scene as in Ref-NeRF [43] has shown improvements over previous state-of-the-art variants, however there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density. It is still common for the NeRF to place regions of density behind surfaces or in front of the camera as detached geometry (floaters) to explain complex phenomena. This leads to a non-realistic geometric scene representation, as in Figure 1.
Applications like robotics and 3D modelling require both accurate geometry and realistic appearance. Poor representation of physical geometry currently poses a large hurdle to the widespread adoption of radiance-based scene representation for geometry critical tasks. Robotic manipulation, autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows. Similarly, traditional structure-from-motion pipelines have great difficulty in reconstructing visually complex objects including reflective and transparent surfaces.
We address geometric inaccuracy arising from view-dependent phenomena with the insight that there are multiple formulations of the plenoptic function which can produce viable scene representations. A surface light field [48] describes the plenoptic function as light originating from a geometrically smooth surface, which we encode using a permutohedral lattice hash encoding. Our insight lies in that we can push the NeRF towards a more geometrically accurate representation in the same rendering framework using additional geometric and appearance based regularisation via curriculum learning. We use a first surface assumption to describe the location of geometry in the scene.
Using a second sampling of the NeRF at these points, we regularise density to produce smoothly-varying normals and thin, continuous sheets of density which more realistically represent scene geometry. Regularising the view-dependent appearance to remove Lambertian components we reduce shape-radiance ambiguity and encouraging more correct structure of the surface. In summary, we make the following contributions:
@@
    We introduce a permutohedral lattice encoding for NeRF which better represents non-planar and complex geometry over traditional cubic hash grids.
  4. 4.
    We introduce the Koala dataset of four scenes with visually challenging objects and complex geometry to benchmark our approach.
We benchmark our approach on state-of-the-art physics based NeRF variants, however our methodology may also be applied to other NeRF frameworks and likely adapted to other radiance based representations.
@@ 2 Related Work
Neural Scene Representations: Implicit neural field approaches [25] produce continuous and high visual fidelity scene representations. These can be queried anywhere within the training set, balancing visual and geometric fidelity. In this work, we improve the geometry of a NeRF and maintain visual fidelity, reducing reliance on introduced geometry to meet scene appearance.
@@
NeRFs [34] leverage the efficiency of ray-based construction similar to light fields with a volumetric scene representation of points which emit light. Subsequent works improved the fidelity of representation [4, 5, 6], recovery of a static scene [40, 42, 33] and performance around view-dependent phenomena such as reflections [43, 23]. However, relying on a purely volumetric rendering approach allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [38, 28, 36] or where light does not follow the physical model of a straight ray through the scene [8]. Additional regularisation terms can dramatically improve the quality of rendering in these cases [36, 28], e.g. improving the accuracy of normals significantly helps representation of complex appearances [32, 43].
More recent works [44] have shown that ray-tracing multiple reflection rays can improve representation reflective scene element, albeit with extremely high computation cost. In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries, maintaining a single-pass volumetric rendering approach.
Radiance Based Scenes: Recent advances in explicit representations like 3D Gaussian splatting [27] (3DGS) have demonstrated improved rendering quality over volumetric implicit models, reflective objects rely heavily on introduced Gaussians behind surfaces to represent view dependence. Although subsequent work has sought to improve the geometric fidelity of the resulting representation [24, 26, 29, 47, 46], these involve substantial changes to the underlying representation or simplified geometry models due to the sparsity of the explicit radiance field. Our approach derives key benefits from the continuity of implicit radiance fields to formulate regularisation terms. We anticipate similarly inspired terms might be formulated for explicit methods like 3DGS, Radiant Foam and others [27, 22, 31, 24].
@@
Surface Representations: Signed distance fields (SDFs) have garnered significant attention [51, 52, 16, 18, 3, 37] as they can represent smooth and continuous geometries in space and may be generated from multi-view constraints alone. For diffuse planar objects, this representation provides a smooth and high-quality reconstruction; however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented [45, 39, 21]. Alternate volumetric encodings better suited to complex geometry like PermutoSDF [39] have substantially improved this representation, which we leverage in this work. Some works have included a reflection parameterisation [21], however, struggle to reconstruct 3D scenes with fine detail, thin structures, and sharp changes compared to volumetric scenes. Appearance and geometry have an intrinsic relationship within neural representations, with geometry being altered to account for non-Lambertian appearance [18]. We maintain a volumetric representation allowing for complex geometry, but with separated view-dependent and -independent appearance.
Appearance Vs. Geometry: Inverse rendering with mesh-based models produces a definite [20] separation of the scene’s appearance, geometry and environment, which are necessary to create realistic renderings of scenes under new conditions. Performing this without prior knowledge of the scene proves to be an immensely difficult task [53, 12, 14, 50], however, utilising physically-based rendering fused with learnt appearance [10, 11] provides sufficient constraints to recover the components of the scene.
@@
Radiance field approaches seeking to learn an accurate scene representation under few view scenarios have used depth consistency [41], additional depth supervision [56] and regularisation using priors [36]. Incorporating an understanding of the physical interactions of light within a scene and its effect on appearance [54] under a solid surface assumption enables accurate geometry to be learnt alongside appearance. In the presence of specularities and other visual phenomena it is difficult to disentangle where appearance has been baked into geometry [49].
Without the need for recovering a bidirectional reflectance distribution function (BRDF), light field approaches [19] provide a more generalised framework. This enables a clear separation between appearance and geometry, reducing the need to acquire environmental view-dependent effects. Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
@@ 3 Method
Figure 2: An overview of the Surf-NeRF methodology. We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene during regular training steps. Sampling at the surface in multiple directions, and at nearby points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region. We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by using a total variation prior. Using neighbouring points we also regularise geometric smoothness on density $\tau$ and consistency of normals $\hat{\mathbf{n}}$ leading to improved, continuous geometry.
@@ 3.1 Preliminaries
Few real surfaces have an entirely view-dependent appearance. Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance and a view-dependent component [48, 30]. A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a piecewise smooth geometry in the scene. This accordingly models phenomena such as volumetric scattering, reflection or transmission as a function of surface position rather than in a volumetric (NeRF) or physics-based (rendering engine) manner. In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance is maintained [48]. Our approach is motivated by this decoupling to encourage Lambertian radiance to exist on smooth sheets of density, or surfaces, with a view-dependent colour.
@@ 3.2 Model
Reflection Parameterisation: Our method builds on the reflection parameterisation in Ref-NeRF [43], which splits the scene into a diffuse and specular appearance term similar to a surface light field. This parameterisation provides enhanced results across Lambertian and specular scenes over original NeRF variants [34, 4] by learning a spatially varying diffuse colour $\mathbf{c}_{d}$, specular tint $\mathbf{s}$, rendering normal $\hat{\mathbf{n}}^{\prime}$ and a view-dependent specular colour $\mathbf{c}_{s}$ for each point in the scene. The final colour of a ray in this parameterisation is given as $\mathbf{c}=\mathbf{c}_{d}+\mathbf{s}\mathbf{c}_{s}$, a linear combination of the diffuse and specular terms. Ref-NeRF struggles to entirely separate Lambertian and specular appearance, utilising density instead to explain complex phenomena such as non-planar reflection, anisotropic reflection and interreflection, leading to the results seen in Figure 1.
Permutohedral Lattice Hash Encoding: We replace the regular cubic hash grid used in prior NeRF variants [7, 35] with a hash-based permutohedral lattice [1]. Prior work [39], has used this structure for its memory efficiency, but our insight lies in the volumetric coverage and relative density of simplices in this structure. A permutohedral lattice, is the most efficient lattice solution the sphere covering problem [15] in low dimensions. Vertices in this lattice are arranged evenly throughout $d$-space and packed more densely [1], allowing for more accurate interpolation of features over cubic volume elements. In $d=3$-space, simplices take the form of regular tetrahedra which can more accurately represent complex non-planar geometries and closely approximating the true geometry of the scene. Our permutohedral lattice encoding therefore has two main benefits over cubic grids: 1) half the memory accesses per interpolation by using barycentric interpolation of a tetrahedron [39], and 2) dramatically improved representation of non-planar surfaces derived from density by representing curved geometry with multi-resolution tetrahedral features.
We apply our regularisation losses denoted by $\mathcal{L}$ and permutohedral lattice encoding to ZipNeRF [6] using the Ref-NeRF parameterisation, leveraging its state of the art performance with grid-based encoding [35]. We make no modifications to the model itself beyond our hash encoding, but include a second, data-driven sampling to impose physically-inspired regularisation. We maintain two proposal networks and one NeRF network with 64 and 32 samples per ray. Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage. An overview of model structure is provided in the supplementary materials.
@@ 3.3 Sampling Radiance at a Surface
We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray intersects a surface in the scene, as shown in Figure 3. These local regularisation terms are based on surface light field properties, current scene geometry and appearance at this location.
@@
We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by the point of highest density closest to the camera. Our candidate surface is the first point along a ray with weight greater than the median weight of the ray, preventing regularisation from occurring behind the true location of the surface. This minimises sampling around points which are occluded and therefore not well positioned with multi-view constraints. Choosing the median ensures that this selection is more robust to skewed weight distributions, particularly early during training. This is shown in Figure 2 left.
Using the point $\mathbf{x}_{\star}$, ray origin $\mathbf{o}$ and direction $\mathbf{d}$, surface normal $\mathbf{\hat{n}}_{\star}$, and covariance $\mathbf{\Sigma}_{\star}$ of this sample’s integrated positional encoding [4], we generate two new batches: a spatial batch to regularise density and a directional batch to regularise appearance. Spatial samples query the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the directional samples query the distribution of view-dependent colour $\mathbf{c}_{s}$ at different viewing angles.
@@
We construct both batches of samples from virtual rays passing through this sphere centred at the surface. The points $p$ define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta. The origins and covariances for these rays,
@@
Using these virtual rays, we construct sampling batches at positions $\mathbf{x}$, and directions $\mathbf{d}$ to employ surface light field regularisation. Both batches inherit covariances from the virtual rays $\mathbf{\Sigma}_{s}$ for anti-aliasing as in ZipNeRF [7].
@@
$\quantity(\mathbf{x}_{\phi},\,\mathbf{d}_{\phi})=\quantity(\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}),$ (2)
@@
where $\sigma_{\star,r}$ is the radial variance of the sample as per mip-NeRF [4]. This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail. Our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training, as shown in the supplement. Since the colours go unused, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.